# Kernel Ridge Regression for conformer ensembles made easy with Structured Orthogonal Random Features

### v1

@@ 4.1 Oxidation potentials in acetonitrile
method MAE (max. $N_{\mathrm{train}}$) est. $N_{\mathrm{train}}$ for target MAE
0.3 eV 0.2 eV
@@
cMBDF $0.19\pm 0.02$ $114$ $446$
SOAP $0.24\pm 0.02$ $156$
FJK $0.28\pm 0.02$ $355$
FJK (s. det.) $0.24\pm 0.03$ $261$
CM $0.35\pm 0.02$
SLATM $0.21\pm 0.03$ $119$
Table 2: Summary of results for machine learning $E_{\mathrm{ox}}$ from Ref. [10], namely MAE for maximum training set size $N_{\mathrm{train}}$ of 474, and values of $N_{\mathrm{train}}$ needed to reach MAE of $0.3$ and $0.2$ eV as estimated from linear interpolation of the log-log plots of MAE vs. $N_{\mathrm{train}}$.
@@ 4.2 Hydration energies
Figure 2 presents MAEs as a function of training set size for all representations considered in this work, with Table 3 summarizing MAEs at the largest training set size (514 datapoints) and estimated training sizes needed to reach chemical accuracy of 1 kcal/mol and FreeSolv’s “default” experimental error [^3] See FreeSolv’s “text notes” data fields. of 0.6 kcal/mol. Relative accuracies of different representations are mostly the same as for LES and are likely explained analogously to the discussion in Subsec. 4.1; the only difference is the pair-determinant FJK being worse than CM in this setup, possibly meaning that the way FJK captures changes of orbitals induced by SMD is disproportionately affected by numerical noise. We also note that introduction of the $\mathcal{E}$ layer (22) did not affect MAE to a degree noticeably above statistical error. Unlike LES, FreeSolv has been used as a benchmark of numerous ML studies, of which several have reported prediction errors of the order of experimental uncertainty [15, 74, 75, 76, 77, 78, 79, 60, 80, 81]. In the context of this work, particularly interesting are the MAEs of 0.417 and 0.359 kcal/mol of Refs. [77] and [79] (which are also the lowest MAEs obtained with a ML algorithm for FreeSolv so far) obtained by first training a model on abundant computational data and then finetuning it [82] for FreeSolv. While similar capitalization on finetuning is impossible in the current MSORF framework, we note a possible avenue to circumvent the problem in Section 5. Our results also improve on errors previously reported for KRR-based approaches: Ref. [12] reported MAE of 0.68 kcal/mol for FML [14] approach combined with FCHL19 representation, while Ref. [60] reported MAE of 1.7 kcal/mol for combination of SOAP descriptors with KRR. Possible reasons for the improvement are discussed in Subsection 4.3.
@@
method MAE (max. $N_{\mathrm{train}}$) est. $N_{\mathrm{train}}$ for target MAE
1.0 kcal/mol 0.6 kcal/mol
[…]
aSLATM (m.-ext.) $0.50\pm 0.04$ $104$ $337$
@@
cMBDF $0.62\pm 0.07$ $106$
cMBDF (m.-ext.) $0.55\pm 0.09$ $132$ $456$
SOAP $0.73\pm 0.04$ $228$
SOAP (m.-ext.) $0.79\pm 0.05$ $259$
FJK $1.53\pm 0.26$
FJK (s. det.) $0.84\pm 0.16$ $234$
FJK (m.-ext.) $1.63\pm 0.29$
FJK (m.-ext., s. det.) $0.91\pm 0.17$ $301$
CM $1.56\pm 0.20$
SLATM $0.59\pm 0.07$ $139$ $499$
@@ 4.3 Comparison of machine learning with ensembles or minimum energy conformers
method MAE (max. $N_{\mathrm{train}}$) est. $N_{\mathrm{train}}$ for target MAE
0.3 eV 0.2 eV
CM $0.36\pm 0.02$
aSLATM $0.19\pm 0.02$ $96$ $440$
FCHL19 $0.17\pm 0.02$ $87$ $316$
@@
method MAE (max. $N_{\mathrm{train}}$) est. $N_{\mathrm{train}}$ for target MAE
1.0 kcal/mol 0.6 kcal/mol
CM $1.72\pm 0.18$
aSLATM $0.51\pm 0.05$ $77$ $350$
FCHL19 $0.49\pm 0.05$ $74$ $301$
@@ 5 Conclusions and outlook
We have devised a general SORF-based procedure for generating physically motivated molecular representations from Boltzmann ensembes of conformers and demonstrated its utility for predicting experimental values for molecular properties. The method both demonstrates systematic decrease of prediction error with increased training set size and allows to reproduce experimental accuracy with hundreds of molecules as training set. At the same time, we did not observe a significant difference between representing a molecule with an ensemble of conformers or a single conformer, implying that, at least for the quantities considered in this work, how conformers are represented is a more important question than how many of them are used.
Prediction errors observed in this work were comparable to state-of-the-art achieved with other ML approaches, despite generating the underlying conformer representations using parameters that (with the exception of SOAP) had not been optimized for the quantities of interest. One possible way to improve on this is to generate conformer representations with another NN, train the combined model on computational data, and then re-optimize the MSORF part of the model on experimental data, essentially finetuning the lowest layers of the combined model and bridging MSORF’s gap with transfer learning techniques, that have shown promise for FreeSolv in Refs. [83, 77, 84, 79]. Another possible direction could be going beyond the simple expressions for weighting contributions of different conformers, atoms, or localized orbitals used in this work to more complicated expressions gleaned from explainable artificial intelligence applied to ab initio calculations [85]. In both cases, it would be possible to capitalize on additional ab initio data while keeping MSORF grounded in KRR formalism. Lastly, while we focused on combining representations of conformers into representations of molecular ensembles, MSORF could analogously combine representations of molecular ensembles into representations of mixtures of molecules. This could make MSORF models useful for exploring chemical spaces of mixtures, a problem relevant for several industrial applications that involve using organic liquids [86, 87, 88].
@@ 6 Code availability
The MSORF algorithm has been made part of the QML2 code [57], with the implementation being based on Numpy [89], Numba [90], and SciPy [91]. The version used in this work together with the ML scripts has been uploaded as the 0.1.5 release at the QML2’s repository. [^4] https://github.com/qml2code/qml2
@@ 7 Supplementary data
Contains data analogous to the one presented in Section 4, but with LES results obtained without shifting $E_{\mathrm{ox}}$ by a constant and FreeSolv results obtained with shifting $E_{\mathrm{hyd}}$ by a stochiometry-dependent term. Also present are text files containing all MAEs for different representations, training set sizes, and random generator seeds.
@@ Appendix A Hyperparameter optimization
Ideally we would want to find hyperparameters that minimize MAE corresponding to $\Delta\vec{q}$, but MAE’s discontinuous derivatives would make it difficult to use gradient during optimization. We thus considered instead using Huber or LogCosh loss functions, both of which can be thought of as “smoothed” versions of MAE with continuous derivatives, but in both cases the smoothness is controlled by parameters that are not straightforward to estimate without knowing beforehand the expected magnitude of model’s errors. We therefore introduced $L_{\mathrm{SCLC}}$, a self-consistent version of LogCosh loss function whose smoothness is controlled by a smoothness parameter $\gamma$, and focused on optimizing
@@
We perform optimization of hyperparameters that can be divided into three groups. The first one are the $\sigma$ hyperparameters mentioned throughout Subsec. 2.2 that affect calculation of $\mathbf{Z}$. The second is $\lambda$ that for fixed $\mathbf{Z}$ affects calculation of $\tilde{\mathbf{K}}$. The third and last, which was mentioned in Section 3, are shifts that can be applied to quantities $\vec{q}$. For the latter, we would attempt to decrease $L_{\mathrm{hyp}}$ by adding to $\vec{q}$ a constant factor of $q^{\mathrm{shift}}$ if the quantity of interest is intensive. If the quantity is extensive, for a point $i$ the quantity $q_{i}$ will be modified by adding
@@
$q^{\mathrm{shift}}_{j}(n_{i})$ is hyperparameter shift specific to nuclear charge $n_{i}$. Note that for fixed $\sigma$ hyperparameters and $\lambda$ $\Delta\vec{q}$ can be written as a linear function of $q^{\mathrm{shift}}$ or $\vec{q}^{\mathrm{shift,tot}}$ that can be evaluated with $\mathcal{O}(N_{\mathrm{train}})$ operations. Linearity in this situation means that $L_{\mathrm{hyp}}$ (28) can be minimized with a quasi-Newton [45] method; in this work the implementation of Limited-memory Broyden–Fletcher–Goldfarb–Shanno algorithm (L-BFGS) [93, 94] in SciPy was used.
@@ Appendix B Self-consistent versions of Huber and LogCosh loss functions
We start our discussion with a general statement about regression loss functions operating on error vectors of dimensionality $N$.
@@
It is straightforward to check that for fixed $\vec{x}$ the second derivative of $L_{\mathrm{LC}}$ w.r.t. $\beta$ is never zero, allowing to calculate the self-consistent version of $L_{\mathrm{LC}}$, $L_{\mathrm{SCLC}}$, numerically via the Newton method as discussed previously.
We implemented both $L_{\mathrm{SCH}}^{2}$ and $L_{\mathrm{SCLC}}^{2}$ in the QML2 code. Though $L_{\mathrm{SCLC}}^{2}$ is significantly more expensive to calculate than $L_{\mathrm{SCH}}^{2}$, we chose the former for our work since its cost is small relative to the cost of finding the SVD decomposition of the $\mathbf{Z}$ matrix (30) and its derivatives are infinitely continuous for all $\vec{x}\neq\vec{0}$ [as can be verified by differentiating (35)], with $\vec{x}=\vec{0}$ extremely unlikely to be encountered during hyperparameter optimization.
As a closing remark, we underline that Equation (35) implies that the $\gamma$ operates similarly to the way $\beta$ operates in the definitions of $L_{\mathrm{H}}$ (46) and $L_{\mathrm{LC}}$ (50). Since the $\beta$ parameter in this context is sometimes described as an absolute error value beyond which a point is considered an outlier, $\gamma$ can be described as how large an error is compared to $L_{\mathrm{SC}}$ (i.e., to the “average smoothed error”) to be considered an outlier. As mentioned in A, we found $\gamma$ to be a more convenient parameter to regulate in our applications.

### v2

@@ 4.1 Oxidation potentials in acetonitrile
method max. $N_{\mathrm{train}}$ est. $N_{\mathrm{train}}$ for target MAE
MAE RMSE 0.3 eV 0.2 eV
@@
[…]
SOAP $0.23\pm 0.01$ $0.33\pm 0.03$ $151$
FJK $0.27\pm 0.04$ $0.36\pm 0.06$ $322$
FJK (s. det.) $0.23\pm 0.02$ $0.34\pm 0.03$ $240$
CM $0.38\pm 0.02$ $0.49\pm 0.03$
SLATM $0.20\pm 0.01$ $0.29\pm 0.04$ $116$
Table 2: Summary of results for machine learning $E_{\mathrm{ox}}$ from Ref. [10], namely MAE and RMSE for maximum training set size $N_{\mathrm{train}}$ of 474, and values of $N_{\mathrm{train}}$ needed to reach MAE of $0.3$ and $0.2$ eV as estimated from linear interpolation of the log-log plots of MAE vs. $N_{\mathrm{train}}$.
@@ 4.2 Hydration energies
Figure 2 presents MAEs as a function of training set size for all representations considered in this work, with Table 3 summarizing MAEs at the largest training set size (514 datapoints) and estimated training sizes needed to reach chemical accuracy of 1 kcal/mol and FreeSolv’s “default” experimental error [^3] See FreeSolv’s “text notes” data fields. of 0.6 kcal/mol. Relative accuracies of different representations are mostly the same as for LES and can be explained analogously to the discussion in Subsec. 4.1; the only difference is pair-determinant FJK being as inaccurate as CM in this setup, possibly meaning that the way FJK captures changes of orbitals induced by SMD is disproportionately affected by numerical noise. We also note that introduction of the $\mathcal{E}$ layer (22) only noticeably affected prediction errors in the case of cMBDF. Unlike LES, FreeSolv has been used as a benchmark in numerous ML studies, of which several have reported prediction errors of the order of experimental uncertainty [15, 76, 77, 78, 79, 80, 81, 60, 82, 83]. In the context of this work, particularly interesting are the MAEs of 0.417 and 0.359 kcal/mol of Refs. [79] and [81] (which are also the lowest MAEs obtained with a ML algorithm for FreeSolv so far) obtained by first training a model on abundant computational data and then finetuning it [84] for FreeSolv. While similar capitalization on finetuning is impossible in the current MSORF framework, we note a possible avenue to circumvent the problem in Section 5. Our results also improve on errors previously reported for KRR-based approaches: Ref. [12] reported MAE of 0.68 kcal/mol for FML [14] approach combined with FCHL19 representation, while Ref. [60] reported MAE of 1.7 kcal/mol for combination of SOAP descriptors with KRR. Possible reasons for the improvement are discussed in Subsection 4.3.
@@
method max. $N_{\mathrm{train}}$ est. $N_{\mathrm{train}}$ for target MAE
MAE RMSE 1.0 kcal/mol 0.6 kcal/mol
@@
[…]
cMBDF $0.63\pm 0.05$ $1.27\pm 0.28$ $100$
cMBDF (m.-ext.) $0.53\pm 0.05$ $0.92\pm 0.12$ $145$ $425$
SOAP $0.76\pm 0.05$ $1.15\pm 0.10$ $228$
SOAP (m.-ext.) $0.75\pm 0.04$ $1.16\pm 0.11$ $232$
FJK $1.47\pm 0.15$ $2.33\pm 0.12$
FJK (s. det.) $0.75\pm 0.06$ $1.27\pm 0.12$ $272$
FJK (m.-ext.) $1.64\pm 0.02$ $2.36\pm 0.09$
FJK (m.-ext., s. det.) $0.80\pm 0.18$ $1.33\pm 0.37$ $306$
CM $1.58\pm 0.17$ $2.37\pm 0.48$
SLATM $0.55\pm 0.05$ $0.95\pm 0.08$ $153$ $466$
@@ 4.3 Comparison of machine learning with ensembles or minimum energy conformers
method max. $N_{\mathrm{train}}$ est. $N_{\mathrm{train}}$ for target MAE
MAE RMSE 0.3 eV 0.2 eV
CM $0.39\pm 0.03$ $0.50\pm 0.04$
aSLATM $0.19\pm 0.01$ $0.28\pm 0.04$ $85$ $383$
FCHL19 $0.17\pm 0.01$ $0.25\pm 0.03$ $74$ $301$
@@
method max. $N_{\mathrm{train}}$ est. $N_{\mathrm{train}}$ for target MAE
MAE RMSE 1.0 kcal/mol 0.6 kcal/mol
CM $1.62\pm 0.20$ $2.41\pm 0.40$
aSLATM $0.50\pm 0.06$ $0.88\pm 0.18$ $77$ $366$
FCHL19 $0.49\pm 0.05$ $0.81\pm 0.14$ $78$ $301$
@@ 5 Conclusions and outlook
We have devised a general SORF-based procedure for generating physically motivated molecular representations from Boltzmann ensembes of conformers and demonstrated its utility for predicting experimental values for molecular properties. The method both demonstrates systematic decrease of prediction error with increased training set size and reaches experimental accuracy with hundreds of molecules as training set. At the same time, we did not observe a significant difference between representing a molecule with an ensemble of conformers or a single conformer, implying that, at least for the quantities considered in this work, how conformers are represented is a more important question than how many of them are used.
Prediction errors observed in this work were comparable to state-of-the-art achieved with other ML approaches, despite generating the underlying conformer representations using parameters that (with the exception of SOAP) had not been optimized for the quantities of interest. One possible way to improve on this is to generate conformer representations with another NN, train the combined model on computational data, and then re-optimize the MSORF part of the model on experimental data, essentially finetuning the lowest layers of the combined model. This would bridge the gap between MSORF and transfer learning techniques, which have shown promise when tested on FreeSolv in Refs. [85, 79, 86, 81]. Another possible direction could be going beyond the simple expressions for weighting contributions of different conformers, atoms, or localized orbitals used in this work to more complicated expressions gleaned from explainable artificial intelligence applied to ab initio calculations [87]. In both cases, it would be possible to capitalize on additional ab initio data while keeping MSORF grounded in KRR formalism. Lastly, while we focused on combining representations of conformers into representations of molecular ensembles, MSORF could analogously combine representations of molecular ensembles into representations of mixtures of molecules. This could make MSORF models useful for exploring chemical spaces of mixtures, a problem relevant for several industrial applications that involve using organic liquids [88, 89, 90].
@@ 6 Code availability
The MSORF algorithm has been made part of the QML2 code [57], with the implementation being based on Numpy [91], Numba [92], and SciPy [93]. The version used in this work together with the ML scripts has been uploaded as the 0.1.6 release at the QML2’s repository. [^4] https://github.com/qml2code/qml2
@@ 7 Supplementary Data
Contains data analogous to the one presented in Section 4, but with LES results obtained with shifting $E_{\mathrm{ox}}$ by a constant and FreeSolv results obtained with shifting $E_{\mathrm{hyd}}$ by a stochiometry-dependent term. Also present are text files containing all MAEs and RMSEs for different representations, training set sizes, and random generator seeds.
@@ Appendix A Hyperparameter optimization
Ideally we would want to find hyperparameters that minimize MAE corresponding to $\Delta\vec{q}$, but MAE’s discontinuous derivatives would make it difficult to use the gradient during optimization. We thus considered instead using Huber or LogCosh loss functions, both of which can be thought of as “smoothed” versions of MAE with continuous derivatives, but in both cases the smoothness is controlled by parameters that are not straightforward to estimate without knowing beforehand the expected magnitude of model’s errors. We therefore introduced $L_{\mathrm{SCLC}}$, a self-consistent version of LogCosh loss function whose smoothness is controlled by a smoothness parameter $\gamma$, and focused on optimizing
@@
We perform optimization of hyperparameters that can be divided into three groups. The first one are the $\sigma$ hyperparameters mentioned throughout Subsec. 2.2 that affect calculation of $\mathbf{Z}$. The second is $\lambda$ that for fixed $\mathbf{Z}$ affects calculation of $\tilde{\mathbf{K}}$. The third and last, which was mentioned in Section 3, are shifts that can be applied to quantities $\vec{q}$. For the latter, we would attempt to decrease $L_{\mathrm{hyp}}$ by adding to $\vec{q}$ a constant factor of $q^{\mathrm{shift}}$ if the quantity of interest is intensive. If the quantity is extensive, for molecule $\mathrm{A}_{i}$ the quantity $q_{i}$ will be modified by adding
@@
$q^{\mathrm{shift}}_{j}(n_{i}^{(\mathrm{A}_{i})})$ is hyperparameter shift specific to nuclear charge $n_{i}^{(\mathrm{A}_{i})}$. Note that for fixed $\sigma$ hyperparameters and $\lambda$, $\Delta\vec{q}$ can be written as a linear function of $q^{\mathrm{shift}}$ or $\vec{q}^{\mathrm{shift,tot}}$ that can be evaluated with $\mathcal{O}(N_{\mathrm{train}})$ operations. Linearity in this situation means that $L_{\mathrm{hyp}}$ (28) can be minimized with a quasi-Newton [95] method; in this work the implementation of Limited-memory Broyden–Fletcher–Goldfarb–Shanno algorithm (L-BFGS) [96, 97] in SciPy was used.
@@ Appendix B Self-consistent versions of Huber and LogCosh loss functions
We start with a general statement about regression loss functions operating on error vectors of dimensionality $N$.
@@
An alternative way to find $L_{\mathrm{SCH}}$ without sorting $|x_{i}|$ is derived by considering the $F_{\mathrm{SC}}$ arising in this case as a function defined as equaling $\tilde{F}_{\mathrm{SC}}^{(j)}$ on each interval of the form $(|x^{\mathrm{s}}_{j}|/(2\gamma),|x^{\mathrm{s}}_{j+1}|/(2\gamma))$, where $\vec{x}^{\mathrm{s}}$ is consists of values of $\vec{x}$ sorted by increasing absolute value and augmented with $x^{\mathrm{s}}_{0}=0$, and $\tilde{F}_{\mathrm{SC}}^{(j)}$ reads
$\tilde{F}_{\mathrm{SC}}^{(j)}(\vec{x},l)=\mathrm{MAE}(\vec{x})-(\gamma+1)l+\frac{1}{N}\sum_{j^{\prime}=1}^{j}\mathrm{max}\left[0,\frac{(x^{\mathrm{s}}_{j^{\prime}})^{2}}{4\gamma l}-|x^{\mathrm{s}}_{j^{\prime}}|+\gamma l\right].$ (50)
For fixed $\vec{x}$ each $\tilde{F}_{\mathrm{SC}}^{(j)}$ is a convex monotonously decreasing function, with $\tilde{F}_{\mathrm{SC}}^{(j+1)}$ differing from $\tilde{F}_{\mathrm{SC}}^{(j)}$ by a nonnegative function. This means each $\tilde{F}_{\mathrm{SC}}^{(j)}$ has only one positive root w.r.t. $l$, which is also smaller than the positive root of $\tilde{F}_{\mathrm{SC}}^{(j+1)}$. Uniqueness of $L_{\mathrm{SCH}}$ also implies that of all the roots of $\tilde{F}_{\mathrm{SC}}^{(j)}$, only one actually lies in the corresponding interval $(|x^{\mathrm{s}}_{j}|/(2\gamma),|x^{\mathrm{s}}_{j+1}|/(2\gamma))$. That means $L_{\mathrm{SCH}}$ can be evaluated similarly to the Newton method, by starting with the initial guess $L_{\mathrm{SCH}}=0$ and calculating each successive guess with Equation (49) with $\sum_{\geq}$ and $\sum_{<}$ evaluated based on the previous guess. The procedure will yield increasing estimates of $L_{\mathrm{SCH}}$ until the exact value is reached, as signified by $\sum_{\geq}$ and $\sum_{<}$ staying the same as in the previous iteration.
@@
It is straightforward to check that for fixed $\vec{x}$ the second derivative of $L_{\mathrm{LC}}$ w.r.t. $\beta$ is never zero, allowing to calculate the self-consistent version of $L_{\mathrm{LC}}$, dubbed $L_{\mathrm{SCLC}}$, numerically via the Newton method as discussed previously.
We implemented both $L_{\mathrm{SCH}}^{2}$ and $L_{\mathrm{SCLC}}^{2}$ in the QML2 code. Though $L_{\mathrm{SCLC}}^{2}$ is significantly more expensive to calculate than $L_{\mathrm{SCH}}^{2}$, we chose the former for our work since its cost is small relative to the cost of finding the SVD of the $\mathbf{Z}$ matrix (30) and its derivatives are infinitely continuous for all $\vec{x}\neq\vec{0}$ [as can be verified by differentiating (35)], with $\vec{x}=\vec{0}$ extremely unlikely to be encountered during hyperparameter optimization. We underline that Equation (35) implies that the $\gamma$ operates similarly to the way $\beta$ operates in the definitions of $L_{\mathrm{H}}$ (46) and $L_{\mathrm{LC}}$ (51). Since the $\beta$ parameter in this context is sometimes described as an absolute error value beyond which a point is considered an outlier, $\gamma$ can be described as how large an error is compared to $L_{\mathrm{SC}}$ (i.e., to the “average smoothed error”) to be considered an outlier. As mentioned in A, we found $\gamma$ to be a more convenient parameter to regulate in our applications.
As a closing remark, we note that $L_{\mathrm{SCH}}^{2}$ and $L_{\mathrm{SCLC}}^{2}$ cannot be straightforwardly used with Stochastic Gradient Descent [103] algorithms since these loss functions are not sums of contributions from different elements of $\vec{x}$. While it should be possible to bypass the issue by using $L_{\mathrm{H}}$ or $L_{\mathrm{SC}}$ and enforcing Equation (35) with a Lagrange multiplier, investigating feasibility and usefulness of such a procedure was beyond the scope of this work.